OpenStudy (anonymous):

Two numbers differ by 10 and have a product of 704. Which of the following equations could be used to find the two numbers? A. x(10 – x) = 704 B. x(10x) = 704 C. x(x – 10) = 704 D. x + (x – 10) = 704

OpenStudy (anonymous):

can anybody explain?

OpenStudy (anonymous):

@nincompoop @whpalmer4 @whydoihavetosignupl @EffImHighOffSkittles @eaileen @R00nnYBraiNsbiG @terenzreignz @uri @IloveYou* @primeralph @prettyboy20202 @ajprincess @AnElephant @SnuggieLad @snorl4x @Directrix @dan815 @divagirl421 @DarkKat @Dmitt @faix95 @GOODMAN @hobbs978 @jhonyy9 @Jack17 @KawasumiKimiko @Luigi0210 @Loujoelou @love_jessika15 @lkkk @Callisto @countonme123 @ChrisJP @nincompoop @mukushla @marylou004 @mlb004 @MissHN182 @thefon19 @ryanjnfish @ladycomical @DDCamp @ryanjnfish @jd1028 @phantomofthemathwork @jtrich13 @ajprincess @PaulaCamacho @laxman03 @beccabee77 @babyslapmafro @whpalmer4 @jtrich13 @resident2013 @sumi29 @R00nnYBraiNsbiG @tfuller @Frostbite @SnuggieLad @raque766 @physicsboffin @Falco276 @e.cociuba @blurbendy @kausarsalley @sakigirl @RANE @Danny_Boy @SomeGirl1999 @Jamierox4ev3r @iforgot @lilkittykat36 @n14r96 @ChanceBBrick @knowel @meldo16 @Helpmeplease99 @Manda_luvsu @baseballer20142 @jhalt @RIUKUS @harpelletjain @Math81 @Frank @WildKat @maurimath @shamim @AliaKnight @MayMay_69 @amyamanda @victoria650 @Social @Whatever1 @xxAshxx @lurrch @djoutten @musicalcat98 @Obviously717 @ihategeometry2425 @Zoe_Bailey @RIUKUS @henendez @shaneetta @niah2411 @harpelletjain @Math81 @WildKat @eidan @Hassan07 @Asuman @Frank @bani2100 @turtlebuddy @maurimath @lizzy89 @dborder @peush @EricaaaaB @yehonatan @tori45 @Ojasmin @Markbruce @april115 @RH @shamim @Ryaan @emilyhaddad @Jamierox4ev3r @RANE @NoelGreco @shkrina @lakenclark @Fazeelayaz @sarahusher @JazzyFrizo16 @music101 @HaperFink22 @whydoihavetosignup1 @domu @ashleyjade @Gikundab @Will.Neko @QuinterrianW_23 @tillman--- @baseballer20142 @Manda_luvsu @Charollete @amyamanda @harpreetsk @Whatever1 @HaydenBurden1996 @xxAshxx @lurrch @djoutten @musicalcat98 @Obviously717 @ihategeometry2425 @shaneetta @DLS @NoelGreco @emilyhaddad @marylou004 @JazzyFrizo16 @music101 @HaperFink22 @Courage. @KingAditya @bradely @Gikundab @Will.Neko @QuinterrianW_23 @mcordray @_anastasia0 @ozhobbits @gabigurl @musicalcat98 @Obviously717 @annemoreno @glamourariss @jbsharpe @WildKat @turtlebuddy @turtlebuddy @peush @yehonatan @tori45 @debrajane @EricaaaaB @linabeanax3 @Ojasmin @Markbruce @confusao @sircrashalot @iizabell @tori45 @yehonatan @peush @maurimath @jbsharpe @glamourariss @annemoreno @Obviously717 @mcordray @_anastasia0 @AppylPye @gabigurl @Imtiaz7 @fringe @khadeeja @jhalt @itsmeeQUENNIE @KingAditya @sambhav__jain @whydoihavetosignup1 @mlb004 @Yasminelove1 @caozeyuan @marylou004 @genius12 @agent0smith @AnElephant @lakenclark @JazzyFrizo16 @HaperFink22 @music101 @danielle02 @HaydenBurden1996 @xxAshxx @Obviously717 @Zoe-bailey @peush @yehonatan @tori45 @debrajane @EricaaaaB @linabeanax3 @Fabien1 @Ojasmin @sircrashalot @ankit042 @caozeyuan

OpenStudy (anonymous):

C would be the Answer I believe: x(x – 10) = 704 (x-10) part gives it away since the two numbers differ by 10. Solving for this you would see \[x(x-10) = 704\] Simplifying becomes\[(x^2-10x) = 704\]Though to solve for that you have to use the method of completing the square. If you do that you find: \[x^{2}-10x+25 = 704\] add 25 to both sides you find\[(x-5)^{2} = 729\] Square root both sides you get (x-5) = 27 or (x-5) = -27 Solve for x you find: x = -22 or x = 32 If you plug those back into the original equation it gives 704.

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

No problem =)